Sadiq is thinking about opening a hardware store. He estimates that it would cost Rs.3,500,000 per year to rent the location and buy the merchandise. In addition, he would have to quit his Rs.800,000 per year job as a financial analyst.

  1. What is Sadiq’s opportunity cost of running the store for a year? 
  2. Sadiq thinks he can sell Rs.4,200,000 worth of hardware in a year. What would his accountant consider the store’s profit? 
  3. How much revenue would the store need to generate for Sadiq to earn zero economic profit? 
1
Expert's answer
2021-06-02T10:47:54-0400

Solution


a.)

opportunity cost= foregone benefits

opportunity cost=Rs. 800,000

b.)

stores profit= TR-implicit cost-explicit cost

implicit cost=foregone benefit=Rs. 800,000

explicit cost=rent=Rs. 3,500,000

TR=4,200,000

profit=4,200,000-3,500,000-800,000= -100,000

profit=Rs.-100,000

c.)

revenue needed

4,200,000+100,000=4,300,000

=Rs.4,300,000